A gas measures \\(1.16 dm^{3}\\) at a temperature of \\(15^{o}C\\) and 309 mmHg. To what celsius temperature must the gas be heated when its pressure rises to 523mmHg and for its volume to remain unchanged?
1
Expert's answer
2019-09-04T04:46:21-0400

Use Gay-Lussac's Law:


"\frac{P_1}{T_1}=\frac{P_2}{T_2}"

P1= 309 mmHg

T1(K) = 15+273 = 288 K

P2= 523 mmHg



"\frac{309}{288}=\frac{523}{T_2}"

T2 (K) = 487 K

T2(0C) = 487-273 = 214 0C